OpenStudy (anonymous):

Max went to the store and bought 22 apples. He bought twice as many red apples as he did green, and he bought one-third as many yellow as he did red. How many of each type of apple did he buy? (Points : 1) 18 red; 8 green; 6 yellow 12 red; 6 green; 4 yellow 10 red; 5 green; 7 yellow 14 red; 7 green; 3 yellow

OpenStudy (anonymous):

r=red g=green y=yellow r=2g y=r3 ⟹ y=2g3 so lets write the equation: r+y+g=22 substitute all the values: 2g+2g3+g=22 find common denominators: 6g3+2g3+3g3=22 add: 11g3=22 multiply by the reciprocal to isolate g: 31111×111g31=222×3111 lets write that a little cleaner: g=2×3 multiply: g=6 now plug that into r=2g and y=2g3: r=2(6) ⟹ r=12 y=2(6)3 ⟹ y=123 ⟹ y=4 now lets test it out: r=12 g=6 y=4 12+4+6=22 ? 22=22 ✓ and thus: Max bought 12 red, 6 green, and 4 yellow apples.

OpenStudy (anonymous):

Max bought 12 red, 6 green, and 4 yellow apples.

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Fiona has a jar filled with 150 coins made up of quarters and nickels. The total value of the coins is $25.10. If x represents the number of quarters in the jar, which equation can be used to find the number of quarters? (Points : 1) 0.25x + 0.05(x – 150) = 25.10 0.05x + 0.25(150 – x) = 25.10 0.05x + 0.25(x – 150) = 25.10 0.25x + 0.05(150 – x) = 25.10

OpenStudy (anonymous):

if we make "x" the quarters, then we can say that x + "nickels amount" = 150 coins and we can also say that "nickels amount" = 150-x solving for "nickels amount" there are 25 cents in 1 quarter, and 5cents in 1 nickel so we can say that, the amount of quarters times 25 cents, "x", plus the "nickels amount" times 5 cents= $25.10 25.10 is really just 2,510 cents or pennies so so we can say that 25x + 5(x-150) = 2,510 "x" + (150-x) = 150

OpenStudy (anonymous):

25x + 5(x-150) = 2,510 #divide both sides by 100 to convert it from pennies to dollars

OpenStudy (anonymous):

so which one is that abc, or d sorry my computer slow

OpenStudy (anonymous):

so whats your choice @BrAnDoN_AAS ?

OpenStudy (anonymous):

@BrAnDoN_AAS

OpenStudy (anonymous):

it can't be A because A is the wrong answer I tried it

OpenStudy (anonymous):

oh sorry I was doing something else for a second

OpenStudy (anonymous):

thats okay

OpenStudy (anonymous):

im having too many issues with connection, one sec

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

D?

OpenStudy (anonymous):

Are you sure its not A

OpenStudy (anonymous):

im very sure it's not

OpenStudy (anonymous):

try c

OpenStudy (anonymous):

im pretty sure its A because heres the link I gott the imformation from http://openstudy.com/study#/updates/52715b1ee4b077aaee83b8f7

OpenStudy (anonymous):

The connection on this site sucks

OpenStudy (anonymous):

^^^

OpenStudy (anonymous):

B. it is 0.05x + 0.25(150 – x) = 25.10 0.05x + 37.5 - 0.25x = 25.10 -0.20x + 37.5= 25.10 -37.5 -37.5 -0.20x = -12.4 x = 62 0.05x + 0.25(150 – x) = 25.10 0.05(62) + 37.5 - 15.5 = 25.10 3.1 + 37.5 - 15.5 = 25.10 25.10 = 25.10 correct!

OpenStudy (anonymous):

did you understand?

OpenStudy (anonymous):

and sorry it took so long, had to do it for all 4 equations, while the connection erased everything lol